Sol Summary – December 30st

Sol: 6
Summary Title: Sol Six: Robot Tricks and a Sediment Mix
Author’s Name: Adriana Brown
Mission Status: Nominal
Sol Activity Status:
Sol 6 began with the crew chipping away at the vat of oatmeal from yesterday morning’s breakfast. Soon after, the first EVA crew of the day, Adi (SCI), Sara (XO), and Nathan (ENG) suited up for a brisk 9:30 AM departure to Cowboy Corner. This EVA was an important one; the objective was to test the performance of E.L.F. As one may recall from earlier reports, E.L.F stands for Electronic LeapFrog. As such, the ability of E.L.F to maneuver over rough terrain was tested. We are happy to report that our little robot friend was able to navigate the rocky terrain, but future tests on jumping and crawling will be required.
45 minutes after the first EVA crew returned, the second EVA crew consisting of Adriana (CO), Eshaana (GEO), Gabe (HSO), and Riya (GHO) set out at noon to a familiar field site: Hab Ridge. At Hab Ridge, two sites were selected to bulk sample a stratigraphic section. The first section was at the base of Hab Ridge where the team, broken into two pairs, collected 45 samples over 75 feet of section. Moving farther up path 1103, the team found a second section to sample, where they collected 48 samples over a 93 foot section. Adriana will use these samples to test weight percent carbonate in the different strata and, back on Earth, run them for carbon isotopes as a proxy for oxygenation in seawater. Their sediment-sampling-efforts (a.k.a. a lot of crawling around in the dirt) were much appreciated by Adriana. Thanks, team!
The thoroughly-chilled EVA crew returned at 2:30 PM to the warm and cozy Hab where their crewmates had prepared some delicious black bean burger nuggets, expertly seasoned mashed potatoes (great job, Adi!), and leftover enchiladas. After a hearty lunch, much of the crew retired for short naps and to their various workstations. As we enjoy our sixth evening on the Red Planet, the crew is looking forward to Sara’s savory tomato basil pancakes and, most of all, sleeping in on our rest day tomorrow!
Look Ahead Plan: No EVAs tomorrow, Sunday will be used as a rest day
Anomalies in work: None
Weather: cold and cloudy with some sparse sun early in the morning
Crew Physical Status: Nominal
EVA: EVA #7 to test robotic capabilities and EVA #8 to sample two measured sections on Hab Ridge
Reports to be filed: Journalist Report, Greenhab Report, Operations Report, Sol Summary, Astronomy Report, Photos, EVA report, Mid-Mission Summary
Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site